# Omics Data Discovery Agents

### 10 Mar 2026 v1

@@ 2.2.2 LLM-Based Information Extraction
Metadata extraction is done in two steps: gathering existing structured metadata such as title, journal, keywords, etc. that are readily available via PMC XML fields, and gathering unstructured metadata from the full text and supplemental. The full article text was processed using a separate LLM (gpt-5) to extract additional keywords and identify the locations of raw data, analysis code, and quantitative data. The purpose of the separate LLM is to limit the possible effects of prompt injections; the external LLM is instructed to return a JSON-formatted response, and the content is processed by functions that do not expose the contents to the agent. The separation can be achieved via programmatic access to models, which is currently readily available from the major providers (e.g. Anthropic’s Client SDKs, Azure’s AI Foundry) or executed locally via Ollama 15. Agents were instructed to provide evidence citations for all extracted values and to explicitly abstain when information could not be reliably determined. Text embeddings were generated using OpenAI’s text-embedding-3-small model and stored in a database for use with semantic similarity search.
@@ 2.4.4 Evaluation of Quantification
Figure 2: Comparison of protein identifications reported by Chen et al. and that obtained by ODDA using the article content to inform parameter selection. Filtering: Both datasets were filtered to remove reverse hits, contaminants, and proteins identified only by site. Proteins were matched by UniProt Majority Protein ID and by gene name. The number of proteins with valid intensity values counted for each of the 6 samples (CCl4-1/2/3, Oil-1/2/3). For each sample, proteins present in both datasets were identified and LFQ intensity values were extracted for shared proteins. Intensities were log10-transformed for correlation analysis, then Pearson correlation was calculated per-sample and overall (pooled across all samples).
@@ 3.3 Agent-Guided Raw Data Quantification
We next evaluated the system’s ability to automatically reprocess raw proteomics data using parameters extracted from article text. We selected the article by Taneera et al. Taneera and others (2026), which describes a DIA proteomics experiment examining the effect of PPP1R1A knockdown on INS-1 cell proteomics.
3.3.1 Parameter Extraction
The agent successfully extracted DIA-NN processing parameters from the article text. Key parameters identified included:
@@ 4.4 Summary
A.H. and J.G.M. conceived the study. A.H. developed the software, conducted experiments, and wrote the manuscript. A.H. and J.G.M. reviewed the manuscript.

### 13 Mar 2026 v2

@@ 2.2.2 LLM-Based Information Extraction
Metadata extraction is done in two steps: gathering existing structured metadata such as title, journal, keywords, etc. that are readily available via PMC XML fields, and gathering unstructured metadata from the full text and supplemental. The full article text was processed using a separate LLM (gpt-5) to extract additional keywords and identify the locations of raw data, analysis code, and quantitative data. The purpose of the separate LLM is to limit the possible effects of prompt injections; the external LLM is instructed to return a JSON-formatted response, and the content is processed by functions that do not expose the contents to the agent. The separation can be achieved via programmatic access to models, which is currently readily available from the major providers (e.g. Anthropic’s Client SDKs, Azure’s AI Foundry) or executed locally via Ollama 16. Agents were instructed to provide evidence citations for all extracted values and to explicitly abstain when information could not be reliably determined. Text embeddings were generated using OpenAI’s text-embedding-3-small model and stored in a database for use with semantic similarity search.
@@ 3.3 Agent-Guided Raw Data Quantification
We next evaluated the system’s ability to automatically reprocess raw proteomics data using parameters extracted from article text. We examine two papers; one from Cheng et al. Cheng and others (2025) describing phosphoproteomics in liver fibrosis, and another by Taneera et al. Taneera and others (2026), which describes a DIA proteomics experiment examining the effect of PPP1R1A knockdown on INS-1 cell proteomics.
3.3.1 Cheng et al.
The article by Cheng et al. Cheng and others (2025) provides its quantitative data as well as its raw data, making it well-suited for comparing results. The raw files were fetched by the agent via ProteomeXchange Vizcaíno et al. (2014), which linked to the files hosted on IProX Ma et al. (2019). The files were then quantified using MaxQuant v2.7.5.0, with the requisite parameter file generated by the agent using MaxQuant’s command line utility and parameter values obtained during article ingestion. After peptide quantification from the raw files, both the article and ODDA datasets were filtered to remove reverse hits, contaminants, and proteins identified only by site. Proteins were matched by UniProt Majority Protein ID and by gene name. The number of proteins with valid intensity values counted for each of the 6 samples (CCl4-1/2/3, Oil-1/2/3). For each sample, proteins present in both datasets were identified and LFQ intensity values were extracted for shared proteins. Intensities were log10-transformed for correlation analysis, then Pearson correlation was calculated per-sample and overall (pooled across all samples). Although the published data identified more proteins in each sample than ODDA, the quantities of the overlapping proteins were highly correlated (r>0.95) - see Figure 2.
Figure 2: Comparison of protein identifications reported by Cheng et al. and those obtained by ODDA using the article content to inform parameter selection.
3.3.2 Taneera et al.
The article from Taneera et al. Taneera and others (2026) included some parameters required for processing, including the quantification method used. Some key parameters include:
@@ 4.4 Summary
A.H. and J.G.M. conceived the study. A.H. developed the software, conducted experiments, and wrote the first draft of the manuscript. J.G.M. obtained funding and supervised the study. A.H. and J.G.M. reviewed the manuscript.